OpenStudy (anonymous):

A die is rolled. The set of equally likely outcomes is {1,2,3,4,5,6}. Find the probability of getting a 3. Express the probability as a fraction reduced to lowest terms.

OpenStudy (kropot72):

If the die is fair, there is an equal probability of rolling each of the six numbers. Let the probability of getting any of the 6 numbers on a single roll of the die = x/6. Then \[6\times \frac{x}{6}=1\] What is the value of x/6 ?

OpenStudy (kropot72):

@Luis_Rivera My reply was intended to show the basis of the answer. The result is the same as you posted: x/6 would easily solve to be 1/6.